OpenStudy (anonymous):

Find the curvature κ(t) of the curve X (t) = ( −4 sint )I + ( −4 sint ) j + ( 3 cost ) K

OpenStudy (anonymous):

I found a similar question and its answer.... QUES:Find the curvature κ(t ) of the curve X(t ) = (3 sin t ) I + (3 sin t ) j + (−5 cos t ) K ANSWER: The curvature of space curve in parametrical form like: X(t) = x(t) · i + y(t) · j + z(t) · k is given by κ(t ) = sqrt[ (z''·y' - y''·z')² + (x''·y' - y''·x')² + (x''·z' - z''·x')² ] / (x'² + y'² + z'²) ^(3/2) ( ' denotes derivative with respect to t) x(t)= 3·sint → x'(t) = 3·cost x''(t) = - 3·sint y(t)= 3·sint → y'(t) = 3·cost y''(t) = - 3·sint z(t)= -5·cost → z'(t) = 5·sint z''(t) = 5·cost The curvature is: κ(t ) = sqrt[ (5·cost·3·cost + 3·sint·5·sint)² + (-3·sint·3·cost + 3·sint·3·cost)² + (- 3·sint·5·sint - 5·cost·3·cost)² ] / (9·cos²t + 9·cos²t + 25·sin²t) ^(3/2) → κ(t ) = sqrt[ (15·cos²t + 15·sin²t)² + (0)² + (- 15·sin²t· - 15·cos²t)² ] / (18·cos²t + 25·sin²t) ^(3/2) With the Pythagorean identity sin²t + cos²t =1 ←→ cos²t = 1- sin²t this simplifies to κ(t ) = sqrt[ (15)² + (- 15)² ] / (18 ·(1 - sin²t) + 25·sin²t) ^(3/2) = 15 · sqrt(2) / (18 + 7·sin²t)^(3/2)

OpenStudy (anonymous):

it said its wrong